# ANote[fnsymbol]

### v3

@@ 3 Speech Synthesis 40M
For the textural corpus (top in Figure 3), we employ the zero-shot in-context generation technique proposed by CosyVoice (Du et al., 2024b) to convert the text into natural-sounding speech. However, many conversational and instruction textural datasets are not ideal for speech synthesis due to the nature of the tasks, which do not facilitate effective speech interaction. To address this issue, we enhance the quality of our synthesised speech data by implementing the following pre-processing rules:
  1. 1.
@@
  2. 2.
    Non-text Element Filtering: texts containing an excessively high proportion of non-text elements are excluded. Specifically, texts with a non-text character ratio exceeding 0.2 are removed. Text characters in this context include punctuation marks (e.g., !, ?), alphabetic characters ([a-z]), and Chinese characters.
  3. 3.
@@
  2. 2.
    Rewriting the question in a conversational tone and generating a conversational response via Qwen2-72B (Chu et al., 2024b).
@@
Finally, leveraging the synthesised audio dataset, we introduce the Nexus-audio testset, which comprises 8.1k Chinese and 2.8k English ASR samples. This testset spans a range of application domains, such as corporate meetings, live broadcasting scenarios, etc.
3.3 Training Strategies
Figure 4: Overview of the training stage in Nexus. The first stage aims to map the speech features into semantic space, the second stage aims to enable the audio instruction-following capability, and the last stage aims to enable speech generation capability.
@@ 3.3 Training Strategies
The first stage aims to facilitate the capacity of LLMs to comprehend and process input speech features. To accomplish this, we utilise 8 million bilingual speech-text pairs for alignment training, covering five training objectives, as illustrated in Table 1. Consistent with the methodology employed in Qwen2.5-VL-7B, our training data is structured in the ChatML format. An example of this format is illustrated as below:
Input Format <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>[audio.wav]<|audio_eos|>Please help me transcribe my current speech into text. <|im_end|> <|im_start|>assistant [text] <|im_end|>
@@
The second stage aims to enable the model to follow speech instructions by incorporating speech-based questions and text-based answers. In this stage, we freeze the visual and audio encoders while unfreezing the visual and audio adapters, and backbone LLM. This strategy is designed to enhance the capability of Nexus to respond accurately to multimodal instructions, thereby improving its performance in understanding and processing both visual and auditory information.
By keeping the core encoder parameters fixed, we ensure that the foundational representations remain stable while allowing the adapters to fine-tune and specialize in the nuances of specific tasks or instructions. This strategic adjustment aims to refine the sensitivity of Nexus to multimodal cues without disrupting the robustness of its underlying architecture.
@@
Different from the previous large speech model connected to the TTS module (Fu et al., 2024c), we exploit a decoder to generate speech end-to-end. First, we train Nexus on the Text-to-Speech dataset to make the model learn text-to-speech alignment. Then, we utilize the speech-to-speech dataset to train the model to learn speech understanding and speech synthesis capabilities. At this stage, only the parameters of the audio decoder are trainable.
4 Evaluation
@@
4.1 Multimodal Tasks Alignment Analysis
In this section, we evaluate the performance of Nexus on the following five tasks: vision understanding, Spoken English QA tasks, ASR, Speech-to-Text translation, and Text-to-Speech tasks. These tasks cover a wide range of modalities and capabilities, providing a comprehensive assessment of the versatility and effectiveness of MLLMs across various domains.
@@ 4.1 Multimodal Tasks Alignment Analysis
We conduct a comprehensive quantitative evaluation of Nexuson a range of vision understanding tasks using eight distinct benchmarks: (1) HallusionBench (Hal) (Guan et al., 2023): Evaluates the model’s susceptibility to hallucination and its handling of visual illusions. (2) MathVista (MathV) (Lu et al., 2023): Assesses mathematical reasoning within visual contexts. (3) OCRBench (OCR) (Fu et al., 2024d): Measures optical character recognition (OCR) capabilities on text-intensive images. (4) Video-MME (Fu et al., 2024b): Examines video question-answering reasoning abilities across six domains over diverse temporal ranges. (5) MMMU (Yue et al., 2024): Challenges the model with multi-disciplinary tasks that demand college-level subject knowledge and deliberate reasoning. (6) AI2D: A dataset comprising over 5,000 grade school science diagrams. (7) MMVet (Yu et al., 2024): Evaluates integrated capabilities including recognition, OCR, domain knowledge, language generation, spatial awareness, and mathematical reasoning. (8) MME (Fu et al., 2024a): Measures both perceptual and cognitive abilities across 14 subtasks.
As illustrated in Table 2, we observe that Nexus achieves strong performance to current popular vision-language MLLMs on the MathV, OCR, and Hal benchmarks. Furthermore, when compared to Qwen2.5-VL, Nexus demonstrates superior performance, highlighting its competitive capabilities in vision understanding and reasoning tasks, indicating our model and training strategy can effectively maintain vision-language alignment capability by incorporating audio modality (Finding1).
Table 2: Evaluation on Vision Understanding Benchmarks. The best two values are shown in bold and underlined. The blue row refers to the main competitors. In MMMU benchmark, * indicates that the measurement is developed over the validation set. We can observe that Nexus shows performance comparable to the leading open-source models and advanced closed-source counterparts.
@@
VILA-1.5 Vicuna-v1.5-13B 44.2 41.1 42.5 39.3 69.9 460.0 45.0 1718.2
LLaVA-Next Yi-34B 51.6 48.8 40.4 34.8 78.9 574.0 50.7 2006.5
CogVLM2 Llama3-Instruct-8B 50.5 42.6 38.6 41.3 73.4 757.0 57.8 1869.5
InternLM-Xcomposer2 InternLM2-7B 56.2 41.4 59.5 41.0 81.2 532.0 46.7 2220.4
Cambrian NousHermes2-Yi-34B 54.2 50.4 50.3 41.6 79.5 591.0 53.2 2049.9
InternVL-Chat-1.5 InternLM2-20B 57.1 46.8 54.7 47.4 80.6 720.0 55.4 2189.6
Ovis1.5 Gemma2-It-9B 58.1 49.7 65.6 48.2 84.5 752.0 53.8 2125.2
InternVL2 InternLM2.5-7B 61.5 51.2 58.3 45.0 83.6 794.0 54.3 2215.1
MiniCPM-V 2.6 Qwen2-7B 57.5 49.8 60.6 48.1 82.1 852.0 60.0 2268.7
Qwen2.5-VL Qwen2.5-7B 56.0 51.8* 61.1 71.7 80.7 877.0 - 2299.1
@@
VITA-1.5-Audio Qwen2-7B - 52.1 66.2 44.9 79.3 732.0 49.6 2352.0
EMova-8B LLaMA-3.1-8B - - 61.1 - 82.8 824.0 55.8 2205.0
Baichuan-Omni-1.5 - 58.2 47.3 51.9 47.8 - - 65.4 2186.9
Megrez-3B-Omni Megrez-3B - 51.8 62.0 50.1 82.0 - - 2315.0
@@
Nexus Qwen2.5-VL-7B 57.0 53.2* 62.1 71.1 81.2 882.0 - 2315.5
@@
Table 3: Evaluation on Audio English QA Benchmarks. The accuracy (%) of different models in English question answering on three sets. The parameter size is derived from the backbone LLMs. Our model achieves top accuracy in the LLaMA Q. benchmark, outperforming the same-period competitor MiniCPM-o2.6-7B.
Model Modality LLaMA Q.$\uparrow$
@@
Mini-Omni-0.5B (Xie & Wu, 2024a) Audio&Text 22.00
@@
Nexus Audio&Text 67.33
@@
In the ASR task, we focus on both Mandarin (CH) and English (EN), evaluating performance on the following benchmarks: AIShell-2 (Du et al., 2018b), Librispeech (Panayotov et al., 2015b), and our real-scenario benchmark. As demonstrated in Table 4, our model achieves the best performance on the real scenario ASR testset, Nexus-audio, indicating the robustness of our model in real life (Finding2).
Table 4: Evaluation on ASR Benchmarks. Nexus has demonstrated strong performance in both real-life Mandarin and English ASR tasks. It outperforms specialized speech models, achieving better results in both languages. AIShell-2 is measured over all categories (Mac/iOS/Android), TC: test-clean, TO: test-other. We can observe that Nexus outperforms other models in real scenarios.
@@
AIShell-2 Nexus-audio Librispeech TC Librispeech TO Nexus-audio
@@
Qwen2-Audio-Instruct-7B 3.00/3.00/2.90 35.45 1.70 4.00 26.12
@@
Mini-Omini2-0.5B(Xie & Wu, 2024a) 117.71 9.80 4.70 54.69
VITA-1.5-7B(Fu et al., 2025) - 127.34 7.20 3.40 122.71
MinMo-7B(Chen et al., 2025b) 4.89/4.76/4.96 - 1.74 3.89 -
@@
As for the S2TT, we exploit CoVoST2 (Wang et al., 2020), which provides massive multilingual S2TT datasets. As shown in Table 5, Nexus demonstrates superior performance compared to specialised speech LLM, Qwen2-Audio-7B-Instruct, in both translation tasks. Nevertheless, our model still falls short of the performance of its contemporary competitor, MiniCPM-o2.6-7B. However, we pre-train our model on a small audio dataset (about 30k hours) at the pre-training stage (Finding3).
Table 5: Evaluation on Speech-to-Text Translation Benchmarks. Nexus demonstrated superior performance compared to Qwen2-Audio-Instruct-7B in both translation tasks, however, it still falls short of the performance of its contemporary competitor due to the relatively modest scale of our pre-training speech dataset (about 30k hours).
@@
Qwen2-Audio-7B (Chu et al., 2024a) 24.40 45.20
Qwen2-Audio-Instruct-7B (Chu et al., 2024a)* 22.90 39.50
MinMo-7B (Chen et al., 2025a)* 25.95 46.68
@@
Finally, we evaluate Nexus on Seed-TTS benchmark Anastassiou et al. (2024) [^5] https://github.com/BytedanceSpeech/seed-tts-eval. As shown in Table 6, our experiments demonstrate that when employing pretrained vocoders (e.g., Fishspeech1.4 or CosyVoice2.0), Nexus achieves performance comparable to its competitors (e.g., CosyVoice1.0 or CosyVoice2.0) (Finding4).
Table 6: Evaluation on Text-to-Speech (TTS) Benchmarks. Nexus demonstrated lower performance in both Mandarin (CH) and English (EN) TTS tasks. Notice: MinMo-7B’s result is based on their manually filtered samples as mentioned in their report Chen et al. (2025a), therefore, we do not compare with it.
@@
MinMo-7B (Chen et al., 2025a) 2.48 2.90
Nexus (Fishspeech1.4) 7.55 13.02
Nexus (CosyVoice2.0) 4.53 4.11
@@
4.2 Representational Alignment Analysis
Unembedding stage
In the unembedding stage, we evaluate representational alignment using ImageNet (Deng et al., 2009), which offers a fine granularity in language-vision alignment, enabling a detailed assessment of representational performance. To get the audio input for each language-vision pair, we apply the pretrained Text-to-Speech LLMs, cosyvoice-300M-SFT checkpoint (Du et al., 2024b). As for the baseline, we compare our model with MLLMs after SFT, including Qwen2.5-VL (Bai et al., 2025), LLaMA3.2-Vision (Meta., 2024), LLaVA family Liu et al. (2023), Phi-3.5-Vision Abdin et al. (2024), and InternVL-8B (Chen et al., 2024).
Table 7: Kernel alignment analysis for unembedding space. The representation for each sample is the averaged token embeddings. The best two values are shown in bold and underlined, we can observe our tri-modal model with audio inputs generally outperforms others on both datasets.
@@ 4.2 Representational Alignment Analysis
Wikipedia Caption: short descriptive sentences
Qwen2.5-VL-7B 0.52148 0.07005 2.09765 0.10194 0.04112 0.19668 0.00173
Llama3.2-Vision-Instruct-11B 0.31347 0.03623 1.29980 0.00968 0.00779 0.22120 0.00050
LLaVA1.5-7B 0.57714 0.02886 1.03808 0.04632 0.01586 0.25615 0.00046
LLaVA1.6-7B 0.57812 0.03935 1.36523 0.07933 0.03998 0.23114 0.00082
Phi3.5-Vision-Instruct 0.04980 0.03027 1.14843 0.01669 0.03890 0.18183 0.00066
InternVL2-8B 0.36914 0.04132 1.55761 0.04732 0.01658 0.21739 0.00093
Nexus 0.48828 0.06990 2.07714 0.08924 0.04101 0.20311 0.00167
vision+audio 0.54000 0.10636 2.79003 0.12403 0.06329 0.25624 0.00230
Figure 5: Averaged kernel-alignment score across different hidden layers. vision+audio: both visual and auditory modalities are concurrently fed into the model. The final fused representation is the last token in the sequence as the comprehensive summary of the integrated modalities. we can observe that incorporating audio modality (green bar) can result in better vision-language alignment at most hidden layers.
@@
5 Conclusion and Future Work
In this work, we propose a comprehensive, industry-level omni-modal LLM pipeline that comprises a modularised end-to-end framework, a lightweight training strategy, and an audio data synthesis pipeline to overcome the current challenges, including limited accessibility to tri-modal datasets, higher computational resource requirements, and the complexity of aligning features across three modalities. To this end, we propose an industry-level omni-perceive and -interactive tri-modal LLM, Nexus.
Extensive experiments validate the usefulness of our proposed pipeline and Nexus. In the visual understanding task, Nexus exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, validating the efficiency of our training strategy. In the Text-to-Speech task, based on pretrained vocoder (e.g., Fishspeech1.4 or CosyVoice2.0), Nexus is comparable to its backbone vocoder on Seed-TTS benchmark. In our real-world ASR testset, Nexus achieves outstanding performance, indicating its robustness in real scenarios.
In the future, we intend to link Nexus with vision generative models under our modularised framework to allow the model to seamlessly generate high-quality, contextually coherent visual outputs. Integrating the Artificial Intelligence-Generated Content (AIGC) capabilities into the current model can further broaden its applicability in creative and industrial domains, such as Embodied Intelligence.
Appendix A Experimental Details
The following tables describe the template we used for each pre-training task:
ASR <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|>Please help me transcribe my current speech into text. <|im_end|> <|im_start|>assistant [text] <|im_end|>
Speech-to-Text Chat <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> <|im_end|> <|im_start|>assistant [response] <|im_end|>
Speech-to-Text Translation <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> What is the English equivalent of this Chinese speech? <|im_end|> <|im_start|>assistant [response] <|im_end|>
Speech Synthesis <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>text<|audio_eos|> <|im_end|> <|im_start|>assistant [audio.wav] <|im_end|>
Speech-to-Speech Chat <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> <|im_end|> <|im_start|>assistant <|audio_bos|>audio.wav<|audio_eos|> <|im_end|>
Automatic Audio Caption <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> please describe the speech. <|im_end|> <|im_start|>assistant [text] <|im_end|>
Speech Emotion Recognition <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user <|audio_bos|>audio.wav<|audio_eos|> what is the emotion in this speech? <|im_end|> <|im_start|>assistant [text] <|im_end|>

### v4

@@ 3 Speech Synthesis 40M 6
For the textural corpus (top in Figure 2), we employ the zero-shot in-context generation technique proposed by CosyVoice (Du et al., 2024a) to convert the text into natural-sounding speech. However, many conversational and instruction textural datasets are not ideal for speech synthesis due to the nature of the tasks, which do not facilitate effective speech interaction. To address this issue, we enhance the quality of our synthesised speech data by implementing the following pre-processing rules:
  1. (1)
@@
  2. (2)
    Non-text Element Filtering: texts containing an excessively high proportion of non-text elements are excluded. Specifically, texts with a non-text character ratio exceeding 0.2 are removed. Text characters in this context include punctuation marks (e.g., !), alphabetic characters ([a-z]), and Chinese characters.
  3. (3)
@@
  2. (2)
    Rewriting the question in a conversational tone and generating a conversational response via Qwen2-72B (Chu et al., 2024a).
@@
Furthermore, we introduce the Nexus-audio testset, which comprises 8.1k Chinese and 2.8k English ASR samples. This testset spans a range of application domains, such as corporate meetings, live broadcasting scenarios, etc.
3.3. Training Strategies
@@
The first stage aims to facilitate the capacity of LLMs to comprehend and process input speech features. To accomplish this, we utilise 8 million bilingual speech-text pairs for alignment training, covering five training objectives, as illustrated in Table 1.
@@
The second stage aims to enable the model to follow speech instructions by incorporating speech-based questions and text-based answers. In this stage, we freeze the visual and audio encoders while unfreezing the visual and audio adapters, and backbone LLM. This strategy is designed to enhance the capability of NEXUS-O to respond accurately to multimodal instructions, thereby improving its performance in understanding and processing both visual and auditory information.
By keeping the core encoder parameters fixed, we ensure that the foundational representations remain stable while allowing the adapters to fine-tune and specialize in the nuances of specific tasks or instructions. This strategic adjustment aims to refine the sensitivity of NEXUS-O to multimodal cues without disrupting the robustness of its underlying architecture.
@@
Different from the previous large speech model connected to the TTS module (Fu et al., 2024c), we exploit a decoder to generate speech end-to-end. First, we train NEXUS-O on the Text-to-Speech dataset to make the model learn text-to-speech alignment. Then, we utilize the speech-to-speech dataset to train the model to learn speech understanding and speech synthesis capabilities. At this stage, only the parameters of the audio decoder are trainable.
4. Evaluation
@@
4.1. Multimodal Tasks Alignment Analysis
In this section, we evaluate the performance of NEXUS-O on the following five tasks: vision understanding, Spoken English QA tasks, ASR, Speech-to-Text translation, and Text-to-Speech tasks. These tasks cover a wide range of modalities and capabilities, providing a comprehensive assessment of the versatility and effectiveness of MLLMs across various domains.
@@ 4.1. Multimodal Tasks Alignment Analysis
We conduct a comprehensive quantitative evaluation of NEXUS-Oon a range of vision understanding tasks using eight distinct benchmarks: (1) HallusionBench (Hal) (Guan et al., 2023): Evaluates the model’s susceptibility to hallucination and its handling of visual illusions. (2) MathVista (MathV) (Lu et al., 2023): Assesses mathematical reasoning within visual contexts. (3) OCRBench (OCR) (Fu et al., 2024d): Measures optical character recognition (OCR) capabilities on text-intensive images. (4) Video-MME (Fu et al., 2024b): Examines video question-answering reasoning abilities across six domains over diverse temporal ranges. (5) MMMU (Yue et al., 2024): Challenges the model with multi-disciplinary tasks that demand college-level subject knowledge and deliberate reasoning. (6) AI2D: A dataset comprising over 5,000 grade school science diagrams. (7) MMVet (Yu et al., 2024): Evaluates integrated capabilities including recognition, OCR, domain knowledge, language generation, spatial awareness, and mathematical reasoning. (8) MME (Fu et al., 2024a): Measures both perceptual and cognitive abilities across 14 subtasks.
Table 2. Evaluation on Vision Understanding Benchmarks. The best two values are shown in bold and underlined. The blue row refers to the main competitors. In MMMU benchmark, * indicates that the measurement is developed over the validation set. We can observe that NEXUS-O shows performance comparable to the leading open-source models and advanced closed-source counterparts.
@@
VILA-1.5(Lin et al., 2024) Vicuna-v1.5-13B 44.2 41.1 42.5 39.3 69.9 460.0 45.0 1718.2
LLaVA-Next(Liu et al., 2024) Yi-34B 51.6 48.8 40.4 34.8 78.9 574.0 50.7 2006.5
InternLM-Xcomposer2(Dong et al., 2024) InternLM2-7B 56.2 41.4 59.5 41.0 81.2 532.0 46.7 2220.4
Cambrian(Tong et al., 2024) NousHermes2-Yi-34B 54.2 50.4 50.3 41.6 79.5 591.0 53.2 2049.9
InternVL-Chat-1.5(Chen et al., 2024a) InternLM2-20B 57.1 46.8 54.7 47.4 80.6 720.0 55.4 2189.6
Ovis1.5(Lu et al., 2024) Gemma2-It-9B 58.1 49.7 65.6 48.2 84.5 752.0 53.8 2125.2
InternVL2(Chen et al., 2024b) InternLM2.5-7B 61.5 51.2 58.3 45.0 83.6 794.0 54.3 2215.1
MiniCPM-V 2.6(Yao et al., 2024) Qwen2-7B 57.5 49.8 60.6 48.1 82.1 852.0 60.0 2268.7
Qwen2.5-VL(Bai et al., 2025) Qwen2.5-7B 56.0 51.8* 61.1 71.7 80.7 877.0 - 2299.1
@@
VITA-1.5-Audio(Fu et al., 2025a) Qwen2-7B - 52.1 66.2 44.9 79.3 732.0 49.6 2352.0
EMova-8B(Chen et al., 2025c) LLaMA-3.1-8B - - 61.1 - 82.8 824.0 55.8 2205.0
Baichuan-Omni-1.5(Li et al., 2024a) - 58.2 47.3 51.9 47.8 - - 65.4 2186.9
Megrez-3B-Omni(Li et al., 2025) Megrez-3B - 51.8 62.0 50.1 82.0 - - 2315.0
@@
NEXUS-O Qwen2.5-VL-7B 57.0 53.2* 62.1 71.1 81.2 882.0 - 2315.5
As illustrated in Table 2, we observe that NEXUS-O achieves strong performance to current popular vision-language MLLMs on the MathV, OCR, and Hal benchmarks. Furthermore, when compared to Qwen2.5-VL, NEXUS-O demonstrates superior performance, highlighting its competitive capabilities in vision understanding and reasoning tasks, indicating our model and training strategy can effectively maintain vision-language alignment capability by incorporating audio modality (Finding1).
@@
Table 3. Evaluation on Audio English QA Benchmarks. The accuracy (%) of different models in English question answering on three sets. The parameter size is derived from the backbone LLMs. Our model achieves top accuracy in the LLaMA Q. benchmark, outperforming the same-period competitor MiniCPM-o2.6-7B.
[…]
@@
Mini-Omni-0.5B(Xie and Wu, 2024b) Audio&Text 22.00
@@
NEXUS-O Audio&Text 67.33
@@
In the ASR task, we focus on both Mandarin (CH) and English (EN), evaluating performance on the following benchmarks: AIShell-2 (Du et al., 2018b), Librispeech (Panayotov et al., 2015b), and our real-scenario benchmark. As demonstrated in Table 4, our model achieves the best performance on the real scenario ASR testset, NEXUS-O-audio, indicating the robustness of our model in real life (Finding2).
Table 4. Evaluation on ASR Benchmarks, where CER and WER refer to Character Error Rate and Word Error Rate, respectively. NEXUS-O has demonstrated strong performance in both real-life Mandarin and English ASR tasks. It outperforms specialized speech models, achieving better results in both languages. AIShell-2 is measured over all categories (Mac/iOS/Android), TC: test-clean, TO: test-other.
@@
AIShell-2 NEXUS-O-audio Librispeech TC Librispeech TO NEXUS-O-audio
@@
Qwen2-Audio-Instruct-7B(Chu et al., 2024b) 3.00/3.00/2.90 35.45 1.70 4.00 26.12
@@
Mini-Omini2-0.5B(Xie and Wu, 2024b) 117.71 9.80 4.70 54.69
VITA-1.5-7B(Fu et al., 2025b) - 127.34 7.20 3.40 122.71
MinMo-7B(Chen et al., 2025a) 4.89/4.76/4.96 - 1.74 3.89 -
@@
As for the S2TT, we exploit CoVoST2 (Wang et al., 2020), which provides massive multilingual S2TT datasets. As shown in Table 5, NEXUS-O demonstrates superior performance compared to specialised speech LLM, Qwen2-Audio-7B-Instruct, in both translation tasks. Nevertheless, our model still falls short of the performance of its contemporary competitor, MiniCPM-o2.6-7B. However, we pre-train our model on a small audio dataset (about 30k hours) at the pre-training stage (Finding3).
Table 5. Evaluation on Speech-to-Text Translation Benchmarks. NEXUS-O demonstrated superior performance compared to Qwen2-Audio-Instruct-7B in both translation tasks, however, it still falls short of the performance of its contemporary competitor due to the relatively modest scale of our pre-training speech dataset (about 30k hours).
@@
Qwen2-Audio-7B(Chu et al., 2024b) 24.40 45.20
Qwen2-Audio-Instruct-7B(Chu et al., 2024b)* 22.90 39.50
MinMo-7B(Chen et al., 2025b)* 25.95 46.68
@@
Finally, we evaluate NEXUS-O on Seed-TTS benchmark (Anastassiou et al., 2024) [^5] https://github.com/BytedanceSpeech/seed-tts-eval. As shown in Table 6, our experiments demonstrate that when employing pretrained vocoders (e.g., Fishspeech1.4 or CosyVoice2.0), NEXUS-O achieves performance comparable to its competitors (e.g., CosyVoice1.0 or CosyVoice2.0) (Finding4).
Table 6. Evaluation on Text-to-Speech (TTS) Benchmarks. NEXUS-O demonstrated lower performance in both Mandarin (CH) and English (EN) TTS tasks. Notice: MinMo-7B’s result is based on their manually filtered samples as mentioned in their report (Chen et al., 2025b), therefore, we do not compare with it.
@@
MinMo-7B(Chen et al., 2025b) 2.48 2.90
NEXUS-O (Fishspeech1.4) 7.55 13.02
NEXUS-O (CosyVoice2.0) 4.53 4.11
@@
4.2. Representational Alignment Analysis
Table 7. Kernel alignment analysis for unembedding space. The representation for each sample is the averaged token embeddings. vision+audio refers to the concatenation between audio and vision embeddings. The best two values are shown in bold and underlined, we can observe that (1) our model without audio inputs can lead to better vision-language alignment than the backbone model on 5 out of 7 metrics, and (2) our model with audio inputs generally outperforms others on 4 out of 7 metrics.
@@ 4.2. Representational Alignment Analysis
Figure 3. Averaged kernel-alignment score across different hidden layers. vision+audio: both visual and auditory modalities are concurrently fed into the model. The final fused representation is the last token in the sequence as the comprehensive summary of the integrated modalities. we can observe that incorporating audio modality (green bar) can result in better vision-language alignment at most hidden layers.
Unembedding stage
In the unembedding stage, we evaluate representational alignment using ImageNet (Deng et al., 2009), which offers a fine granularity in language-vision alignment, enabling a detailed assessment of representational performance. To get the audio input for each language-vision pair, we apply the pretrained Text-to-Speech LLMs, cosyvoice-300M-SFT checkpoint (Du et al., 2024a). As for the baseline, we compare our model with MLLMs after SFT, including Qwen2.5-VL (Bai et al., 2025), LLaMA3.2-Vision (Meta., 2024), LLaVA family (Liu et al., 2023), Phi-3.5-Vision (Abdin et al., 2024), and InternVL-8B (Chen et al., 2024b).
@@
5. Conclusion and Future Work
In this work, we propose a comprehensive, industry-level omni-modal LLM pipeline that comprises a modularised end-to-end framework, a lightweight training strategy, and an audio data synthesis pipeline to overcome the current challenges, including limited accessibility to tri-modal datasets, higher computational resource requirements, and the complexity of aligning features across three modalities. To this end, we propose an industry-level omni-perceive and -interactive tri-modal LLM, NEXUS-O.
Extensive experiments validate the usefulness of our proposed pipeline and NEXUS-O. In the visual understanding task, NEXUS-O exhibits superior performance compared with its backbone model - Qwen2.5-VL-7B, validating the efficiency of our training strategy. In the Text-to-Speech task, based on pretrained vocoder (e.g., Fishspeech1.4 or CosyVoice2.0), NEXUS-O is comparable to its backbone vocoder on Seed-TTS benchmark. In our real-world ASR testset, NEXUS-O achieves outstanding performance, indicating its robustness in real scenarios. In the future, we intend to link NEXUS-O with vision generative models under our modularised framework to allow the model to seamlessly generate high-quality, contextually coherent visual outputs. Integrating the Artificial Intelligence-Generated Content (AIGC) capabilities into the current model can further broaden its applicability in creative and industrial domains, such as Embodied Intelligence.